OpenStudy (anonymous):

Given \(\mathbb{F}\) is an ordered field, prove 1>0.

OpenStudy (anonymous):

That's where I don't quite understand. How can you use some wrong assumption to prove something right?

geerky42 (geerky42):

Maybe to prove a contradiction I guess we assumed something wrong to be true then we point out contradiction, hence original statement (1>0) is true.

geerky42 (geerky42):

"if 1<0, then -1>0, then so is (-1)(-1) = 1 by Positive scaling. But 1 is not positive (allegedly). This is a contradiction. So 1 is positive after all. QED"

geerky42 (geerky42):

Sh*t man, I don't ever know what ordered field is... What class are you taking?

geerky42 (geerky42):

@RolyPoly

OpenStudy (anonymous):

Introduction to Math. Analysis

geerky42 (geerky42):

aha, it's far away from where I am now, I think. Does my explanation help somehow?

OpenStudy (anonymous):

I understand that we are trying to prove it by contradiction but still, I am not able to comprehend it.

geerky42 (geerky42):

yeah. me too ha... Sorry, at least I tried.

geerky42 (geerky42):

Maybe @ganeshie8 @agent0smith can help?

OpenStudy (anonymous):

It's ok. Thanks for your help!

OpenStudy (anonymous):

Ordered fields have two important axioms (along with the usual field axioms): \[(1)~~~~\text{If }a,b,c\in\mathbb{F}\text{ with }a\le b,\text{ then }a+c\le b+c\\ (2)~~~~\text{If }0\le a\text{ and }0\le b,\text{ then }0\le ab\] The contradiction suggested by geerky uses the second axiom to prove there is a contradiction.

OpenStudy (anonymous):

Hmm... What I am learning is a bit different, but essentially the same. For (2), \(\text{if } a,b > 0\text{, then }a+b>0\text{ and }a*b >0\)

OpenStudy (anonymous):

I don't know what's wrong with me, but seems my problem is solved. Thanks everyone! Proposition 1: If F is an ordered field and \(a \in \mathbb{F}\) with \(a\ne0\), precisely one of a>0 or -a<0 holds [proof omitted] Suppose 1<0, by proposition 1, we have -1>0 By O2, we have (-1) * (-1) > 0 And by some tedious proof, we have (-1) * (-1) = 1 Hence, we have 1>0, which contradicts to our assumption. Since \(1\ne0\), hence, we have 1>0.